# Evaluating LLMs at Detecting Errors in LLM Responses

### v1

@@ Abstract
With Large Language Models (LLMs) being widely used across various tasks, detecting errors in their responses is increasingly crucial. However, little research has been conducted on error detection of LLM responses. Collecting error annotations on LLM responses is challenging due to the subjective nature of many NLP tasks, and thus previous research focuses on tasks of little practical value (e.g., word sorting) or limited error types (e.g., faithfulness in summarization). This work introduces ReaLMistake, the first error detection benchmark consisting of objective, realistic, and diverse errors made by LLMs. ReaLMistake contains three challenging and meaningful tasks that introduce objectively assessable errors in four categories (reasoning correctness, instruction-following, context-faithfulness, and parameterized knowledge), eliciting naturally observed and diverse errors in responses of GPT-4 and Llama 2 70B annotated by experts. We use ReaLMistake to evaluate error detectors based on 12 LLMs. Our findings show: 1) Top LLMs like GPT-4 and Claude 3 detect errors made by LLMs at very low recall, and all LLM-based error detectors perform much worse than humans. 2) Explanations by LLM-based error detectors lack reliability. 3) LLMs-based error detection is sensitive to small changes in prompts but remains challenging to improve. 4) Popular approaches to improving LLMs, including self-consistency and majority vote, do not improve the error detection performance. Our benchmark and code are provided at https://github.com/psunlpgroup/ReaLMistake.
Figure 1: Left: Tasks in existing LLM evaluation benchmarks are often subjective and not suitable for collecting errors made by LLMs for the purpose of evaluating binary error detection methods. Right: We introduce ReaLMistake benchmark with realistic, objective, and diverse errors made by LLMs for evaluating error detection.
@@
Benchmark Task Response Models # Data Label Involve Subjective Criteria Error Categories
@@
Table 1: Comparison between ReaLMistake and prior benchmarks for evaluating error detectors or evaluators for LLM responses. ReaLMistake is the first benchmark that includes objective and realistic errors in LLM responses in diverse categories.
@@ 1 Introduction
An obstacle in studying error detection is the lack of benchmarks that include binary error annotations (i.e., whether the response contains errors or not) on objective, realistic, and diverse errors made by LLMs. An NLP task should satisfy the following criteria to be suitable for collecting errors made by LLMs for the purpose of benchmarking error detection. First, to provide objective error labels, tasks should not involve subjectivity or ambiguity. As illustrated in Figure 1, subjective tasks on which humans cannot provide reliable error annotation are not suitable for benchmarking error detection methods. For example, in text summarization, humans cannot objectively annotate binary error labels because the evaluation of “whether the generated summary properly includes important content in the source” (the “relevance” criterion) is subjective (Laban et al., 2023). We cannot evaluate error detectors if humans cannot determine whether LLM responses include errors. As shown in Table 1, tasks in benchmarks for ranking-based LLM evaluators (Zheng et al., 2023; Wang et al., 2023a; 2024) and criterion-specific error detectors (e.g., factuality evaluators) (Laban et al., 2023; Manakul et al., 2023) often involve subjectivity (Zeng et al., 2024). Second, error detection benchmarks should be built on tasks that make LLMs introduce errors similar to errors in real-world applications. Some prior studies use tasks such as word sorting (Tyen et al., 2024) and chess (Wu et al., 2023), which are objective but have different properties from practical uses of LLMs. Third, to provide a comprehensive testbed, the benchmarks need to include a variety of tasks to provide errors in diverse categories. Previous research in detecting mistakes in LLM responses focuses on errors of limited types (Laban et al., 2023; Manakul et al., 2023; Lightman et al., 2024), as shown in Table 1. Finally, tasks should be challenging for strong LLMs to make them introduce errors.
@@ 2 Related Work
Ranking-based Evaluation
Evaluation of LLMs has been mainly studied in ranking-based (pairwise comparison) (Chen et al., 2023c; Dubois et al., 2023; Zheng et al., 2023; Wang et al., 2023a; 2024; Zeng et al., 2024) or Likert scale (Chiang & Lee, 2023) since recent LLMs often do not make obvious mistakes and many NLP tasks involve subjectivity. These evaluation approaches are useful for system-level comparison of LLMs but not suitable for detecting erroneous responses. Binary error detection is motivated by a real-world situation in which we want to automatically detect errors in LLM responses to reject or improve bad outputs.
Self-Improvement
Error detection is an important task in its own right, but it has been mainly studied in the context of improving LLM outputs by using feedback from LLMs (Bai et al., 2022; Madaan et al., 2023; Shinn et al., 2023; Kim et al., 2023; Gou et al., 2023; Chen et al., 2023a; Huang et al., 2024). These studies often evaluate the performance of the refined outputs on downstream tasks and an analysis of the binary error detection step is lacking.
@@ 3.1 How to Collect Objective, Diverse, and Realistic Errors in LLM Responses?
(1) Create Tasks that Make LLMs Introduce Objective, Realistic, and Diverse Errors
@@
(2) Create Tasks Challenging for LLMs but Feasible for Humans
@@
To address these problems, we design tasks involving properties that are observed to be challenging for LLMs: (a) following requirements (Liu et al., 2023c; Zhou et al., 2023a), (b) comparison of multiple texts (Laban et al., 2023; Chen et al., 2023b), and (c) detection of small mistakes in text (Cohen et al., 2023). As shown in Table 2, tasks in our benchmark make GPT-4 introduce errors in more than 50% of cases, while their inputs consist of a few hundred words that only involve high-school-level mathematics and knowledge in Wikipedia. These tasks allow us to efficiently collect errors made by LLMs.
@@ 3.2 Tasks
Our benchmark includes three challenging, objective, and realistic tasks designed for collecting diverse categories of errors in LLM responses. As discussed in the previous section, these tasks are designed so that they can be evaluated by the four objective and diverse criteria, eliminating the subjectivity in the error annotation process while ensuring the diversity. In addition, the three tasks are selected to involve skills required in real-world applications of LLMs (e.g., arithmetic reasoning, fact verification, and factual knowledge), in contrast to the objective tasks in prior work that have properties different from those of real-world uses of LLMs, such as chess (Wu et al., 2023) and word sorting (Tyen et al., 2024). Figure 2 includes an example from each task. Figure 3 visualizes the creation processes.
@@ 1.0 Pro
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
@@ 5 Conclusion
We provide a GitHub repository [^5] https://github.com/psunlpgroup/ReaLMistake that includes the ReaLMistake benchmark (also hosted at Hugging Face Datasets), [^6] https://huggingface.co/datasets/ryokamoi/realmistake all outputs from LLM-based error detectors used in Section 4, and evaluation code.
@@ Acknowledgment
We thank Greg Durrett for the valuable discussions and comments. This work was supported by a Cisco Research Grant.
@@ A.3 Human-written Instruction of Evaluation Steps
Error Detector Gemma Llama 2 Mistral Qwen 1.5
@@ B.3 Answerability Classification
Factual Mistakes in Numbers
@@
Other Factual Mistakes
@@
Time Sensitive Questions
@@ Appendix D Diversity of Tasks in ReaLMistake
Math Word Problem Generation
@@
Fine-grained Fact Verification
@@
Answerability Classification Task
@@ G.1 Additional Results for Comparison to Other Tasks
Table 10 includes LMSYS Chatbot Arena Elo Rating and MMLU performance (five-shot) of 12 LLMs used in Figure 5 and Figure 9. We accessed LMSYS Chatbot Arena on March 16, 2024.888https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard Sources of MMLU performances are provided in the table.
Refer to captionRefer to caption

### v2

@@ Abstract
With Large Language Models (LLMs) being widely used across various tasks, detecting errors in their responses is increasingly crucial. However, little research has been conducted on error detection of LLM responses. Collecting error annotations on LLM responses is challenging due to the subjective nature of many NLP tasks, and thus previous research focuses on tasks of little practical value (e.g., word sorting) or limited error types (e.g., faithfulness in summarization). This work introduces ReaLMistake, the first error detection benchmark consisting of objective, realistic, and diverse errors made by LLMs. ReaLMistake contains three challenging and meaningful tasks that introduce objectively assessable errors in four categories (reasoning correctness, instruction-following, context-faithfulness, and parameterized knowledge), eliciting naturally observed and diverse errors in responses of GPT-4 and Llama 2 70B annotated by experts. We use ReaLMistake to evaluate error detectors based on 12 LLMs. Our findings show: 1) Top LLMs like GPT-4 and Claude 3 detect errors made by LLMs at very low recall, and all LLM-based error detectors perform much worse than humans. 2) Explanations by LLM-based error detectors lack reliability. 3) LLM-based error detection is sensitive to small changes in prompts but remains challenging to improve. 4) Popular approaches to improving LLMs, including self-consistency and majority vote, do not improve the error detection performance. Our benchmark and code are provided at https://github.com/psunlpgroup/ReaLMistake.
Figure 1: Left: Tasks in existing LLM evaluation benchmarks are often subjective and not suitable for collecting errors made by LLMs for the purpose of evaluating binary error detection methods. Right: We introduce the ReaLMistake benchmark with realistic, objective, and diverse errors made by LLMs for evaluating error detection.
@@
Benchmark Task Response Models # Data Label No Subjective Criteria Error Categories
@@
Table 1: Comparison between ReaLMistake and prior benchmarks for evaluating error detectors or evaluators for LLM responses. There is a deficiency in benchmarks for evaluating pointwise evaluation, although there are diverse benchmarks for evaluating pairwise evaluation (e.g., Zheng et al., 2023; Dubois et al., 2023; Zhang et al., 2023). ReaLMistake is the first benchmark that includes objective and realistic errors in LLM responses in diverse categories.
@@ 1 Introduction
An obstacle in studying error detection is the lack of benchmarks that include binary error annotations (i.e., whether the response contains errors or not) on objective, realistic, and diverse errors made by LLMs. [^1] In this paper, “objective errors” represent errors in LLM responses that do not involve ambiguity or subjectivity in the detection process and can be detected in a binary manner with a high inter-annotator agreement. An NLP task should satisfy the following criteria to be suitable for collecting errors made by LLMs for the purpose of benchmarking error detection. First, to provide objective error labels, tasks should not involve subjectivity or ambiguity. As illustrated in Figure 1, subjective tasks on which humans cannot provide reliable error annotation are not suitable for benchmarking error detection methods. For example, in text summarization, humans cannot objectively annotate binary error labels because the evaluation of “whether the generated summary properly includes important content in the source” (the “relevance” criterion) is subjective (Laban et al., 2023). It is difficult to evaluate error detectors if humans cannot determine whether LLM responses include errors. As shown in Table 1, tasks in benchmarks for ranking-based LLM evaluators (Zheng et al., 2023; Wang et al., 2023a; 2024) and criterion-specific error detectors (e.g., factuality evaluators) (Laban et al., 2023; Manakul et al., 2023) often involve subjectivity (Zeng et al., 2024). Second, error detection benchmarks should be built on tasks that make LLMs introduce errors similar to errors in real-world applications. Some prior studies use tasks such as word sorting (Tyen et al., 2024) and chess (Wu et al., 2024), which are objective but have different properties from practical uses of LLMs. Third, to provide a comprehensive testbed, the benchmarks need to include a variety of tasks to provide errors in diverse categories. Previous research in detecting mistakes in LLM responses focuses on errors of limited types (Laban et al., 2023; Manakul et al., 2023; Lightman et al., 2024), as shown in Table 1. Finally, tasks should be challenging for strong LLMs to make them introduce errors.
@@ 2 Related Work
Ranking-based Evaluation.
Evaluation of LLMs has been mainly studied in ranking-based (Chen et al., 2023c; Dubois et al., 2023; Zheng et al., 2023; Wang et al., 2023a; 2024; Zeng et al., 2024) or Likert scale (Chiang & Lee, 2023) since recent LLMs often do not make obvious mistakes and many NLP tasks involve subjectivity. These evaluation approaches are useful for system-level comparison of LLMs but not suitable for detecting erroneous responses. Binary error detection is motivated by a real-world situation in which we want to automatically detect errors in LLM responses to reject or improve bad outputs.
Self-Correction.
Error detection is an important task in its own right, but it has been mainly studied in the context of improving LLM responses using feedback from LLMs (Bai et al., 2022; Madaan et al., 2023; Shinn et al., 2023; Kim et al., 2023; Chen et al., 2023a; Gou et al., 2024; Pan et al., 2024; Kamoi et al., 2024). These studies often evaluate the performance of the refined outputs on downstream tasks and an analysis of the binary error detection step is lacking.
@@ 3.1 How to Collect Objective, Diverse, and Realistic Errors in LLM Responses?
(1) Create Tasks that Make LLMs Introduce Objective, Realistic, and Diverse Errors.
@@
(2) Create Tasks Challenging for LLMs but Feasible for Humans.
@@
To address these problems, we design tasks involving properties that are observed to be challenging for LLMs: (a) following requirements (Liu et al., 2024; Zhou et al., 2023a), (b) comparison of multiple texts (Laban et al., 2023; Chen et al., 2023b), and (c) detection of small mistakes in text (Cohen et al., 2023). As shown in Table 2, tasks in our benchmark make GPT-4 introduce errors in more than 50% of cases, while their inputs consist of a few hundred words that only involve high-school-level mathematics and knowledge in Wikipedia. These tasks allow us to efficiently collect errors made by LLMs.
@@ 3.2 Tasks
Our benchmark includes three challenging, objective, and realistic tasks designed for collecting diverse categories of errors in LLM responses. As discussed in the previous section, these tasks are designed so that they can be evaluated by the four objective and diverse criteria, eliminating the subjectivity in the error annotation process while ensuring the diversity. In addition, the three tasks are selected to involve skills required in real-world applications of LLMs (e.g., arithmetic reasoning, fact verification, and factual knowledge), in contrast to the objective tasks in prior work that have properties different from those of real-world uses of LLMs, such as chess (Wu et al., 2024) and word sorting (Tyen et al., 2024). Figure 2 includes an example from each task. Figure 3 visualizes the creation processes.
@@ 4.4 Popular Techniques to Improve LLMs Do Not Improve LLM-based Error Detection
5 Future Directions - Evaluation on Ambiguous Errors
In this work, we propose to evaluate error detectors on objective errors because this approach has multiple advantages, including smaller noises, the absence of subjectivity of annotators, binary annotations, and smaller annotation costs. However, since many NLP tasks are subjective, we expect future research to evaluate whether error detectors have a good correlation with human evaluators on ambiguous errors. A possible approach to collecting benchmarks for evaluating error detectors on ambiguous errors is to collect annotations by multiple annotators for each instance to get average error scores. However, difficulties in creating such benchmarks include (1) hiring annotators who represent the distribution of (possibly subjective) error annotations by general people, (2) evaluating binary predictions from LLM-based error detectors using labels by multiple annotators, and (3) high annotation cost. We expect future work to address these challenges.
@@ 6 Conclusion
Experiments.
We provide a GitHub repository [^6] https://github.com/psunlpgroup/ReaLMistake that includes the ReaLMistake benchmark (also hosted at Hugging Face Datasets [^7] https://huggingface.co/datasets/ryokamoi/realmistake), all outputs from LLM-based error detectors used in Section 4, and evaluation code.
Annotation.
Appendix B provides annotation instructions we used for our benchmark.
@@ Acknowledgment
This work was supported by a Cisco Research Grant. We thank Greg Durrett for the valuable discussions and comments. We appreciate valuable suggestions from anonymous reviewers.
@@ A.3 Human-written Instruction of Evaluation Steps
Error Detector
Gemma
Llama 2 Mistral Qwen 1.5
@@ B.3 Answerability Classification
Factual Mistakes in Numbers.
@@
Other Factual Mistakes.
@@
Time Sensitive Questions.
@@ Appendix D Diversity of Tasks in ReaLMistake
Math Word Problem Generation.
@@
Fine-grained Fact Verification.
@@
Answerability Classification Task.
@@ G.1 Additional Results for Comparison to Other Tasks
Table 10 includes LMSYS Chatbot Arena Elo Rating and MMLU performance (five-shot) of 12 LLMs used in Figure 5 and Figure 9. We accessed LMSYS Chatbot Arena on March 16, 2024.999https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard Sources of MMLU performances are provided in the table.